OpenStudy (anonymous):

sin^-1(-.18) = -.18 correct?

OpenStudy (bahrom7893):

No.. why? I mean i don't know but in general, Sin^-1(x) does not equal x

OpenStudy (anonymous):

unlikely

OpenStudy (turingtest):

\[\sin^{-1}x=\theta\]tells you what angle theta gives you the value sin(theta)=x example\[\sin(\frac\pi2) =0\to\frac\pi2=\sin^{-1}0\]

OpenStudy (anonymous):

I'm just punching this in the calculator and that is answer I got. I was just startled that x and the answer were the samee

OpenStudy (turingtest):

\[\sin x=x\]only has one solution: x=0